OpenStudy (anonymous):

Use the quadratic formula to solve 3x2 – 12 = 0.

OpenStudy (saifoo.khan):

2 or -2

OpenStudy (anonymous):

\[3(x^2-4)=0\] \[3(x-2)(x+2)=0\] \[x-2=0\] \[x=2\] \[x+2=0\] \[x=-2\] there!